# A Review of Design Concerns in Superconducting Quantum Circuits

### v2

@@ I Introduction
Superconducting circuits are a leading quantum computing technology, combining strong couplings and long-lived coherence with flexible engineering and scalability[1]. By carefully arranging of capacitors, inductors, transmission lines, and Josephson junctions, experimenters create circuits with coherent quantized energy levels, often called artificial atoms. While every atom of a particular isotope is completely identical, artificial atoms are extremely customizable. Their behavior is determined by the circuit graph, which shows the arrangement of each circuit element and the and connections between them; the circuit element parameters, i.e., the exact values of the inductances, capacitances, impedances, antenna couplings, etc.; the layout, i.e., the physical geometry of superconducting metal, insulating substrate, dielectric layers, etc.; the embedding structure, i.e., the physical enclosure that houses a circuit and routes signals to it; and the materials used in fabricating the circuit as well as the fabrication processes that affect surface and interface layers. Each of these aspects must be carefully engineered to create the desired behavior.
@@
The review is structured according to the typical workflow of a design loop. It begins with a brief description of the physics of superconducting quantum circuits at an effective Hamiltonian level. We then discuss depictions of these circuits as being composed of lumped elements in a particular circuit graph, and how these depictions can be used to design the desired effective Hamiltonian. We follow this by discussing how a physical device layout is designed and how the layout can effect behavior. We then discuss simulations that can be used to predict device behavior given a layout. We describe how the device can change between layout and fabrication. And finally we close the loop by describing how measured device behavior can be used to improve the next round of designs. Throughout each discussion we describe how the steps interact with each other. The goal of this review is not to provide a comprehensive, detailed description of each aspect of design—in many cases these articles have already been written, and we refer the reader to them throughout the text. Instead this review is intended as an overview that can be used as an introduction to superconducting device design for new researchers in the field, and a centralized reference that they can use to find more detailed resources on each aspect of the problem.
@@ II Superconducting Qubits
Here we give a brief introduction to the behavior of superconducting circuits. More comprehensive reviews can be found in references [2, 3, 4, 5, 1, 6, 7, 8].
Figure 1: Diagram of an inductor (left, orange), Josephson junction (center, red), and capacitor (blue, right), forming a parallel LJC circuit. Each element has an associated energy. Tiling LJC subcircuits in series and parallel allows one to build up an arbitrary circuit.
The fundamental building block of a superconducting circuit is the parallel LJC subcircuit, with an inductor (with inductance $L$), Josephson junction (with critical current $I_{0}$), and capacitor (with capacitance $C$) in parallel. See Fig. 1 for a diagram. This produces the Hamiltonian
@@
Here $\hat{n}$ is the charge number operator, $n_{g}$ is a constant that accounts for the effect of voltage bias across the capacitor, $\hat{\phi}$ is the dimensionless gauge-invariant node flux (that is, the phase difference across the junction/inductor–see Section III.1), and $\phi_{ext}$ is the dimensionless external flux threading the junction-inductor loop $\phi_{ext}=\frac{\Phi_{ext}}{2\varphi_{0}}$, where $\varphi_{0}\equiv\hbar/2e$ is the reduced magnetic flux quantum. The behavior of this subcircuit can be entirely captured by the component energies $E_{c}\equiv e^{2}/2C$, $E_{J}\equiv\varphi_{0}I_{0}$, and $E_{L}\equiv\varphi_{0}^{2}/L$, along with the environmental bias variables $n_{g}$ and $\phi_{ext}$. Note that a time-dependent $\phi_{ext}$ can lead to another term, but we will ignore this for simplicity [9, 10, 11].
Varying the ratios of the energies produces different qubit types, including the ubiquitous transmon ($E_{L}=0$, $E_{J}\gg E_{c}$) and the many varieties of fluxonium ($E_{c}\sim E_{J}\gg E_{L}$), as well as harmonic oscillators ($E_{J}=0$). Subcircuits such as this can be placed in series or parallel to generate an arbitrary lumped element circuit—omission of an element can be represented as setting its energy to 0.
This circuit Hamiltonian description is accurate and, as we will see, is the level at which device designers must work. Unfortunately it is often quite inconvenient to work with when designing quantum architectures, algorithms, and gates, and for most quantum information science theory. Instead, the circuit Hamiltonian is typically reduced to an effective Hamiltonian description, treating the circuit as a set of modes of oscillation which are best described as qubits, qudits, harmonic oscillators, and weakly anharmonic oscillators. These modes are coupled, typically with couplings that are weak enough that the original modes still form a useful basis. As an illustrative example, consider a single transmon coupled to a linear $LC$ resonator via a coupling capacitor as in Fig. 2. The circuit-level Hamiltonian is
@@
A common simplifying assumption is to treat the transmon as a qubit and drop the non-excitation-conserving couplings $ab$ and $a^{\dagger}b^{\dagger}$ (i.e., take the rotating wave approximation), in which case one recovers the well-studied Jaynes-Cummings Hamiltonian [13]:
@@
This depiction is conceptually simple, but is often not accurate enough for, e.g., designing gate schemes [14, 15, 16], understanding measurement-induced state transitions [17, 18], or interpreting spectroscopic results [19, 20] (see below). One message we hope the reader takes away from this review is: do not neglect the higher transmon levels or the non-rotating-wave-approximation terms unless you are certain they will not affect the result!
@@ III Circuit Graph and Parameters
Figure 4: (a) Scanning electron microscope image of a superconducting device. False color highlights indicate two CPW resonators (red and teal), a tunable coupler (blue), a fixed-frqeuency transmon (purple), and coupling capacitors (yellow). The blue line at the top is a transmission line throguh, while the unhighlighted line to the left is a fast flux line inductively coupled to the coupler’s SQUID. (b) Intermediate depiction of the device, showing circuit elements and their arrangement. (c) Circuit graph of the device, with nodes indicated.
@@ III.1 From Graph to Hamiltonian
As described in Section II, we can quantize a circuit by writing its Hamiltonian in terms of charge and flux operators defined at each node. It is therefore required to transform from branch coordinates to node operators. This can be a complex procedure, which is well described in [27, 28] (an alternative approach using loop fluxes is described in [29]); here we give a brief summary.
@@
We then define a spanning tree of capacitances, which connects every node to ground through a path of only capacitive couplings without forming any loops. This means that every node has a single unique path to ground through a series of capacitive branches. Summing the branch fluxes from ground to the $i$th mode thus defines the node fluxes $\phi_{i}$; any external bias flux threading a loop is taken to add on branches which are not part of the tree. We can then use node fluxes as our coordinates and write the potential energy in terms of these coordinates. For instance, an inductor $L_{ij}$ connecting nodes $i$ and $j$ will contribute energy $E=(\phi_{i}-\phi_{j}+\phi_{ext,ij})^{2}/2L_{ij}$, where $\phi_{ext,ij}$ is the bias flux assigned to this branch of the loop. Likewise, the energy contribution from a Josephson junction depends on the difference in node fluxes across it (plus any bias flux assigned to that part of a loop).
@@ IV.1 Circuit Parameters
The first priority in layout design is to create a device with the correct circuit parameters–capacitances, inductances, transmission line resonances, distributed couplings, and Josephson energies. Josephson energy is the most straightforward to design for—one simply designs a Josephson junction with a junction area that gives the correct critical current, given the critical current density of a fabrication process. Edge effects in junctions may mean that the critical current does not precisely scale with area for small junctions, but these can generally be accounted for phenomenologically once a fabrication process is well-characterized.
@@ IV.2 Crosstalk
Figure 5: Circuit diagram of a typical multi-path coupling between two transmon qubits. Combining a direct capacitive coupling with an intermediate coupling resonator (purple) allows the designer to cancel the effective $ZZ$ interaction induced by the transmons’ higher excited states.
@@ IV.3 Connectivity
Another consideration is the capacitance budget of a subcircuit. Each capacitive coupling between a subcircuit and its neighbors adds in parallel to the subcircuit capacitance(s), which must be reduced to compensate and maintain a constant total. Eventually the coupling capacitances dominate the total capacitance, and it becomes impossible to increase a coupling without weakening all others. This is especially a concern in subcircuits such as fluxonium qubits, where the charge coupling matrix elements are quite weak [64] and so coupling capacitances must be large [65, 66]. A similar problem can occur with inductive couplings. One alternative is to use galvanic coupling, where two subcircuits share a linear or nonlinear inductive element (i.e., an inductor or Josephson element). By using a SQUID as the galvanic coupler, it is possible to tune the coupling rapidly in-situ [67, 68], and the galvanic coupling budget is generally quite large.
These concerns exist in the context of quantum error correction (QEC) architecture design. In general, the more connected the qubit graph, the easier QEC is to accomplish. That is, if each qubit can be made to run high-fidelity two-qubit gates with a greater number of other qubits, then fault-tolerant QEC can be achieved with fewer resources (qubits and operations) and at a higher gate error rate [69]. At some point it becomes impractical to add more fixed couplings. Directly coupling two distant qubits together then requires some long-range controllable interaction such as controlled transmission of microwave photons [70, 71, 72], resonant couplings between modular elements [73], or measurement-based remote entanglement [74, 75, 76].
@@ IV.4 Interface Losses
It has now been well established that interface layers between substrate and metal (superconductor), between substrate and vacuum, and between metal and vacuum can all act as dielectrics that are many orders of magnitude more lossy than crystalline silicon/sapphire dielectric substrates [77, 78]. The decay rate $\Gamma$ of a qubit as limited by dielectric losses is
@@
where $\omega_{q}$ is the qubit frequency, $P_{i}\in[0,1]$ is the participation ratio of a dielectric material (the fraction of total electric field energy that is in this material), $\tan\delta_{i}$ is the microwave loss tangent of that material, and $\beta$ is a dimensionless prefactor indicating the strength of the charge decay matrix element. For a harmonic oscillator $\beta=1$ and for a transmon $\beta\approx 1$.
Figure 6: Diagram of the electric field distribution between two capacitor plates that are close together (a) or far apart (b). For the same total field energy, the close-together plates concentrate more field in surface and interface layers (purple, red, blue), which may be much lossier than the substrate (green) or vacuum.
@@
Another method to reduce interface losses is to design the geometric features with smaller perimeters, fewer sharp corners, and fewer narrow sections, all of which can increase the participation of lossy surface layers [82]. Similarly, metal films incorporating oxide layers (for junction fabrication) may be lossier than single-layer films [83]. These films are typically only present in the connection between Josephson elements and larger-scale features; reducing the lengths of these connectors can reduce the loss.
@@ IV.5 Radiation
A circuit may couple to the external electromagnetic field via antenna modes. Devices that are meant to be embedded in 3D cavities are typically designed with large antenna structures to provide dipole coupling to one or more cavity modes. This coupling can be modeled several ways, as discussed in Section V.5. Typical design rules for microwave antennae apply, including matching the antenna length to the wavelength (or some integer divisor), loading the ends with capacitive structures, and avoiding spurious antenna resonances at unwanted frequencies.

### v3

@@ I Introduction
Superconducting circuits are a leading quantum computing technology, combining strong couplings and long-lived coherence with flexible engineering and scalability [1]. By carefully arranging of capacitors, inductors, transmission lines, and Josephson junctions, experimenters create circuits with coherent quantized energy levels, often called artificial atoms. While every atom of a particular isotope is completely identical, artificial atoms are extremely customizable. Their behavior is determined by the circuit graph, which shows the arrangement of each circuit element and the connections between them; the circuit element parameters, i.e., the exact values of the inductances, capacitances, impedances, antenna couplings, etc.; the layout, i.e., the physical geometry of superconducting metal, insulating substrate, dielectric layers, etc.; the embedding structure, i.e., the physical enclosure that houses a circuit and routes signals to it; and the materials used in fabricating the circuit as well as the fabrication processes that affect surface and interface layers. Each of these aspects must be carefully engineered to create the desired behavior.
@@
A typical design process may be summarized in steps:
  1. 1.
    Determine the desired behavior, i.e., the desired Hamiltonian.
  2. 2.
    Find a circuit model that implements this Hamiltonian.
  3. 3.
    Create a physical device layout that the designer guesses will implement the circuit model.
  4. 4.
    Perform finite-element electromagnetic simulations of the layout and construct a circuit model and/or Hamiltonian based on the simulation results.
  5. 5.
    Compare the model/Hamiltonian from the simulation to the desired Hamiltonian. If they do not match to within the required tolerance, alter the layout, simulate, compare, and iterate steps 3-5 until the predicted behavior matches the desired behavior.
  6. 6.
    Fabricate the device, measure its behavior, and compare to the simulation output (which matches target behavior). If experiment results deviate significantly from simulation, refine the simulation and/or fabrication pipelines and repeat steps 3-6.
The review is structured according to this design loop workflow. It begins with a brief description of the physics of superconducting quantum circuits at an effective Hamiltonian level. We then discuss depictions of these circuits as being composed of lumped elements in a particular circuit graph, and how these depictions can be used to design the desired effective Hamiltonian. We follow this by discussing how a physical device layout is designed and how the layout can effect behavior. We then discuss simulations that can be used to predict device behavior given a layout. We describe how the device can change between layout and fabrication. And finally we close the loop by describing how measured device behavior can be used to improve the next round of designs. Throughout each discussion we describe how the steps interact with each other. The goal of this review is not to provide a comprehensive, detailed description of each aspect of design—in many cases these articles have already been written, and we refer the reader to them throughout the text. Instead this review is intended as an overview that can be used as an introduction to superconducting device design for new researchers in the field, and a centralized reference that they can use to find more detailed resources on each aspect of the problem. LABEL:tab:references summarizes these references, organized by topic.
@@ II Superconducting Qubits
Here we give a brief introduction to the behavior of superconducting circuits. More comprehensive reviews can be found in references [2, 3, 4, 5, 1, 6, 7, 8, 9].
Figure 1: Diagram of an inductor (left, orange), Josephson junction (center, blue), and capacitor (teal, right), forming a parallel LJC circuit. Each element has an associated energy. Tiling LJC subcircuits in series and parallel allows one to build up an arbitrary circuit.
A circuit can be modeled as a set of nodes (each with its own voltage) connected by branches (each carrying its own current). A branch can be formed by any lumped circuit element. Neglecting loss, in a superconducting circuit such an element can be an inductor (with inductance $L$), a Josephson junction (with critical current $I_{0}$), or a capacitor (with capacitance $C$) (we neglect distributed elements for the moment). One can combine these into the parallel LJC subcircuit. See Fig. 1 for a diagram. Note that the capacitance here may be entirely due to the Josephson junction geometry or may have some additional contribution from a physical shunting capacitor, and the inductance is a physical shunting inductance in parallel with the junction. This is the simplest possible circuit that captures all possible branches between two nodes. It produces the Hamiltonian
@@
Here $\hat{n}$ is the charge number operator (where charge is quantized in units of Cooper pairs, i.e., $2e$), $n_{g}$ is a constant that accounts for the effect of voltage bias across the capacitor, $\hat{\phi}$ is the dimensionless gauge-invariant node flux (that is, the phase difference across the junction/inductor–see Section III.1 for a rigorous definition), and $\phi_{ext}$ is the dimensionless external flux threading the junction-inductor loop $\phi_{ext}=\frac{\Phi_{ext}}{\varphi_{0}}$, where $\varphi_{0}\equiv\hbar/2e$ is the reduced magnetic flux quantum. Note that node flux is distinct from flux threading a loop, as explained in Section III.1. The behavior of this subcircuit can be entirely captured by the component energies $E_{c}\equiv e^{2}/2C$, $E_{J}\equiv\varphi_{0}I_{0}$, and $E_{L}\equiv\varphi_{0}^{2}/L$, along with the environmental bias variables $n_{g}$ and $\phi_{ext}$. A time-dependent $\phi_{ext}$ can lead to another term, but we will ignore this for simplicity [10, 11, 12].
While the parallel LJC subcircuit need not be a qubit, one can produce certain qubit types from a single LJC circuit by varying the energies. This includes the ubiquitous transmon ($E_{L}=0$, $E_{J}\gg E_{c}$) and the many varieties of fluxonium ($E_{c}\sim E_{J}\gg E_{L}$), as well as harmonic oscillators ($E_{J}=0$). Subcircuits such as this can be placed in series or parallel to generate an arbitrary lumped element circuit—omission of an element can be represented as setting its energy to 0. Recent work has enumerated all possible combinations of such 2-node building blocks into circuits with up to 5 nodes [13], providing pre-solved Hamiltonians for such circuits. General circuit Hamiltonians can be built up as described in Section III.1. We have omitted loss (i.e., resistance) from this picture, which we will discuss later in Section IV.4.
This circuit Hamiltonian description is accurate and, as we will see, is the level at which device designers must work. Unfortunately it is often inconvenient to work with when designing quantum architectures, algorithms, and gates, and for most quantum information science theory. Instead, the circuit Hamiltonian is typically reduced to an effective Hamiltonian description, treating the circuit as a set of modes of oscillation which are best described as qubits, qudits, harmonic oscillators, and weakly anharmonic oscillators. These modes are coupled, typically with couplings that are weak enough that the original modes still form a useful basis. As an illustrative example, consider a single transmon coupled to a linear $LC$ resonator via a coupling capacitor as in Fig. 2. The circuit-level Hamiltonian is
@@
A common simplifying assumption is to treat the transmon as a qubit and drop the non-excitation-conserving couplings $ab$ and $a^{\dagger}b^{\dagger}$ (i.e., take the rotating wave approximation, RWA), in which case one recovers the well-studied Jaynes-Cummings Hamiltonian [15]:
@@
This depiction is conceptually simple, but is often not accurate enough for, e.g., designing gate schemes [16, 17, 18], understanding measurement-induced state transitions [19, 20], or interpreting spectroscopic results [21, 22] (see below). One message we hope the reader takes away from this review is: do not neglect the higher transmon levels or the non-RWA terms unless you have already confirmed they will not affect the result! This means that one must first calculate behavior including higher levels and non-RWA terms and compare to the behavior without. Then, if there is no significant difference, one could neglect the higher levels in any situation where the qubit approximation / RWA is expected to work even better. For
ratio of anharmonicity to other energy scales (e.g. drive strength) is even higher, but not situations where it is lower! Likewise, non-RWA terms could be neglected if the ratio of their oscillation frequencies to other frequencies in the problem were even larger, but should be kept otherwise.
@@ III Circuit Graph and Parameters
Figure 4: (a) Scanning electron microscope image of a superconducting device. False color highlights indicate two CPW resonators (purple and lilac), a tunable coupler (red), a fixed-frequency transmon (orange), and coupling capacitors (teal). The blue line at the top is a transmission line through, while the unhighlighted line to the left is a fast flux line inductively coupled to the coupler’s SQUID. (b) Intermediate depiction of the device, showing circuit elements and their arrangement, with colors matching the coloring in (a). Distributed elements are shown as transmission lines. Full subcircuits, including the schematic lines connecting different elements to a node, are colored to show identification of subcircuits with their own resonant modes (c) Circuit graph of the device, with nodes indicated. Ports are changed to lumped-element resistors and distributed elements are replaced with effective lumped models. Elements are colored according to their subcircuits in (b), but connections within nodes are left black to indicate no pre-identification of modes
@@ III.1 From Graph to Hamiltonian
As described in Section II, we can quantize a circuit by writing its Hamiltonian in terms of charge and flux operators defined at each node. It is therefore required to transform from branch coordinates to node operators. This can be a complex procedure, which is well described in refs. [29, 30] (an alternative approach using loop fluxes is described in [31]); here we give a brief summary. An excellent review with simple worked examples can be found in ref. [9].
@@
We then define a spanning tree of capacitances, which connects every node to ground through a path of only capacitive couplings without forming any loops. This means that every node has a single unique path to ground through a series of capacitive branches. Summing the branch fluxes from ground to the $i$th node thus defines the node fluxes $\Phi_{i}$, which we divide by the reduced flux quantum $\varphi_{0}\equiv\hbar/2e$ to get the dimensionless node flux. Any external bias flux threading a loop is taken to add on branches which are not part of the tree. We can then use node fluxes as our coordinates and write the potential energy in terms of these coordinates. For instance, an inductor $L_{ij}$ connecting nodes $i$ and $j$ will contribute energy $E=(\phi_{i}-\phi_{j}+\phi_{ext,ij})^{2}/2L_{ij}$, where $\phi_{ext,ij}$ is the dimensionless bias flux assigned to this branch of the loop. Likewise, the energy contribution from a Josephson junction depends on the difference in node fluxes across it (plus any bias flux assigned to that part of a loop).
@@ IV.1 Circuit Parameters
The first priority in layout design is to create a device with the correct circuit parameters–capacitances, inductances, transmission line resonances, distributed couplings, and Josephson energies. Josephson energy is the most straightforward to design for—one simply designs a Josephson junction with a junction area that gives the correct critical current, given the critical current density of a fabrication process. Edge effects in junctions may mean that the critical current does not precisely scale with area for small junctions (empirically, less than $\sim$ 200 nm on a side), but these can generally be accounted for phenomenologically once a fabrication process is well-characterized.
@@ IV.2 Crosstalk
Figure 5: Circuit diagram of a typical multi-path coupling between two transmon qubits (within the dashed orange boxes). Combining a direct capacitive coupling with an intermediate coupling resonator (purple) allows the designer to cancel the effective $ZZ$ interaction induced by the transmons’ higher excited states.
@@ IV.3 Connectivity
Another consideration is the capacitance budget of a subcircuit. Each capacitive coupling between a subcircuit and its neighbors adds in parallel to the subcircuit capacitance(s), which must be reduced to compensate and maintain a constant total. Eventually the coupling capacitances dominate the total capacitance, and it becomes impossible to increase a coupling without weakening all others. This is especially a concern in subcircuits such as fluxonium qubits, where the charge coupling matrix elements are quite weak [66] and so coupling capacitances must be a significant fraction of the overall capacitance [67, 68]. A similar problem can occur with inductive couplings. One alternative is to use galvanic coupling, where two subcircuits share a linear or nonlinear inductive element (i.e., an inductor or Josephson element). By using a SQUID as the galvanic coupler, it is possible to tune the coupling rapidly in-situ [69, 70], and the galvanic coupling budget is generally large enough to allow for coupling to at least 4 other qubits (although complexity may make this approach undesirable).
These concerns exist in the context of quantum error correction (QEC) architecture design. In general, the more connected the qubit graph, the easier QEC is to accomplish. That is, if each qubit can be made to run high-fidelity two-qubit gates with a greater number of other qubits, then fault-tolerant QEC can be achieved with fewer resources (qubits and operations) and at a higher gate error rate [71]. At some point it becomes impractical to add more fixed couplings. Directly coupling two distant qubits together then requires some long-range controllable interaction such as controlled transmission of microwave photons [72, 73, 74, 75], resonant couplings between modular elements [76], or measurement-based remote entanglement [77, 78, 79].
@@ IV.4 Interface Losses
It has now been well established that interface layers between metal (superconductor) and substrate, between substrate and air (or vacuum), and between metal and air (or vacuum) can all act as dielectrics that are many orders of magnitude more lossy than crystalline silicon/sapphire dielectric substrates [80, 81]. Microwave loss tangents of substrates are of order $\tan\delta\sim 10^{-6}-10^{-7}$ while interfaces can be as lossy as $\tan\delta\sim 10^{-2}-10^{-3}$. The decay rate $\Gamma$ of a qubit as limited by dielectric losses is
@@
where $\omega_{q}$ is the qubit frequency, $P_{i}\in[0,1]$ is the participation ratio of the ith dielectric material (the fraction of total electric field energy that is in this material), $\tan\delta_{i}$ is the microwave loss tangent of that material, and $\beta$ is a dimensionless prefactor indicating the strength of the charge decay matrix element. For a harmonic oscillator $\beta=1$ and for a transmon $\beta\approx 1$.
Figure 6: Qualitative depiction of the electric field distribution between two capacitor plates that are close together (a) or far apart (b). For the same total field energy, the close-together plates concentrate more field in surface and interface layers. The metal-air (MA) interface is shown in purple, the metal-substrate (MS) interface in red, and the substrate-air (SA) in blue; interfaces are drawn as rough boundaries to emphasize their amorphous nature. These interfaces are typically much lossier than the substrate (green) or vacuum, with microwave loss tangents 3 to 5 orders of magnitude higher than bulk silicon or sapphire.
@@
Another method to reduce interface losses is to design the geometric features with smaller perimeters, fewer sharp corners, and fewer narrow sections. All of these changes reduce the participation of lossy surface layers [85]. Similarly, metal films incorporating oxide layers (for junction fabrication) may be lossier than single-layer films [86]. These films are typically only present in the connection between Josephson elements and larger-scale features; reducing the lengths of these connections can reduce the loss.
@@ IV.5 Radiation
A circuit may couple to the external electromagnetic field via antenna modes. Devices that are meant to be embedded in 3D cavities are typically designed with large (mm-scale) antenna structures to provide dipole coupling to one or more cavity modes. This coupling can be modeled several ways, as discussed in Section V.5. Typical design rules for microwave antennae apply, including matching the antenna length to the wavelength (or some integer divisor), loading the ends with capacitive structures, and avoiding spurious antenna resonances at unwanted frequencies.